OpenStudy (anonymous):

What is the simplest form of the quotient? 3^sqrt x3/5^sqrt x2

OpenStudy (anonymous):

\[\frac{ \sqrt[3]{x^3} }{ \sqrt[5]{x^2} }\]

OpenStudy (anonymous):

\[\sqrt[3]{x} = x ^{\frac{ 1 }{ 3 }}\] for the numerator we have: \[(x^{3})^{\frac{ 1 }{ 3 }} = x ^{\frac{ 3 }{ 3 }} = 1\] Now you try for the denominator